The common stock of KRJ Enterprises is currently selling for $38.76 per share. Last year the company reported net income of $359 million. Included on its income statement was total depreciation expense of $54 million. There were 200 million shares of common stock outstanding. Using the proxy for cash flow, what is the company's price/cash flow ratio? Present your answer to two decimal places. e.g. 20.00.
Answer #1

Cash flow =Depreciation+Net income =359+54 =413 million

Cash flow per share =413/200 =2.065

Price/cash flow ratio =38.76/2.065 =18.77